OpenStudy (anonymous):

Help Please!!!! I'll Fan and give medal!!! Find the common solution for 2x + 7y = -38 and -5x - 8y = 38 X=? Y=?

OpenStudy (texaschic101):

use the addition/elimination method...

OpenStudy (anonymous):

2x+7y = -38 -5x-8y=38 the thing is to make the coefficient of one of the variables equal and then eliminate it. multiplying the first equation by 5, 10x+35y = -190 second equation by 2, -10x-16y=76 adding them eliminates the x-variable 35y - 16y = -190 + 76 Therefore, 19y = -114 find y from here, then substitute in any one of the originally given equations and get the values of x.

OpenStudy (anonymous):

y=-6 andx=2

OpenStudy (anonymous):

VERY GOOD :)